OpenStudy (anonymous):

can some1 plz do out this integral step by step 4 me. 2/x^2+4x+3 and integrate between 2 and 0. I have it done but i think i may have done it wrong as i got a minus on the denominator and so it doesnt fit into the tan-1 formula

OpenStudy (amistre64):

factor the denom decomp it all then integrate up into ln parts

OpenStudy (amistre64):

(x+1)(x+3) 2 = A/(x+1) + B/(x+3) determine A and B when x=-1 and x=-3

OpenStudy (amistre64):

then it should just integrate to: Aln(x+1) + Bln(x+3) ; from 0 to 2

OpenStudy (anonymous):

Thank you. I forgot about the partial fraction technique!

OpenStudy (amistre64):

i always tend to forget about one technique or another as well :)